Vinny’s Overhead Construction had free cash flow during 2015 of $36.1 million. The change in gross fixed assets on Vinny’s balance sheet during 2015 was $3.0 million and the change in net operating working capital was $4.0 million.

Calculate the missing amounts on Vinny’s income statement below. (Enter your answers in millions of dollars rounded to 2 decimal places. Input all amounts as positive values.)

VINNYS OVERHEAD CONSTRUCTION, CORP Income Statement for Year Ending December 31, 2015 (in millions of dollars) Net sales Less: Cost of goods sold 198.10 124.1 Gross profits Less: Other operating expenses 74.00 20.40 Earnings before interest, taxes, depreciation, and amortization (EBITDA) Less: Depreciation 53.60 18.60 35 Earnings before interest and taxes (EBIT) Less: Interest Earnings before taxes (EBT) Less: Taxes Net income 15.55
Answer #1

Solution:

free cash flow = EBIT (1- tax) + Depreciation - capex - change in working capital

36.1 = 35 ( 1- tax) + 18.6 - 3 - 4

36.1 - 11.6 = 35 - 35 * tax

35 * tax = 10.5

tax = 10.5/35 = .3 = 30%

Let interest be x then Earning before tax = 35-x

Tax = .3*(35-x)

EBT - tax = net income

35 -x - .3(35-x) = 15.55

.7(35-x) = 15.55

24.5 - .7x = 15.55

.7 * x = 24.5-15.55 = 8.95

x = 8.95/.7 = 12.78

Tax = .3 *(35-12.78) = 6.66

EBIT 35
less: Interest 12.78
Earning before tax 22.22
Tax 6.66
Net Income 15.55